What is the answer to (p-1)2=6 ?
Answer by Alan3354(69443)

What is the answer to (p-1)2=6 ?
-------------
It depends on the question.
If the question is "what's the degree of the polynomial?", the answer is 1.
If the question is "Is the equation linear?", the answer is yes.
-------
If you want to solve for p, you should say, "solve for p."
(p-1)2=6
p-1 = 3
p = 4